# Ultracoherent superconducting cavity-based multiqudit platform with error-resilient control

### v2

@@ Abstract
Bosonic quantum systems have emerged as a compelling platform for quantum computing, offering several intrinsic advantages over traditional qubit-based architectures. By encoding quantum information in harmonic oscillator modes, these systems benefit from long coherence times, large Hilbert spaces, and relatively simple error channels dominated by photon loss. Coupling to nonlinear superconducting circuits has enabled universal control and gate operations using a variety of techniques [1, 2, 3, 4, 5, 6, 7, 8] in both single-mode [9] and multimode [10, 11] circuit QED systems implemented with three-dimensional (3D) SRF cavities featuring millisecond-scale coherence times. These capabilities have enabled hardware-efficient implementations of bosonic quantum error correction, including demonstrations that surpass the break-even point [12, 13, 14]. More recently, these control strategies have begun to extend to ultra-high-coherence cavities [15, 16], with coherence times exceeding tens of milliseconds and the potential for further improvements [17], opening new avenues for scalable quantum information processing.
@@
Qubit-based quantum computing has long been the dominant paradigm. In bosonic error correction, cavity modes are typically used to encode logical qubits [18, 19]. The large Hilbert space of a cavity also makes it a natural platform for implementing a qudit using multiple quantum levels, both natively and with error correction [20]. Qudits offer enhanced information density and can enable more compact and efficient quantum circuits across a wide range of algorithms [21, 22, 23]. They are particularly well-suited for quantum simulations in chemistry [24, 25], condensed matter [26], and high-energy physics (HEP) [27, 28], where they can represent physical degrees of freedom such as large spins, rotors, and gauge fields [29]. In HEP applications, a quantum field can also be discretized in position space and encoded within a single high-dimensional qudit [30]. The emergence of high-coherence cavities presents a promising path for improved control over high-photon-number Fock states, thereby enabling qudit-based quantum computing and simulation.
@@
Despite their advantages, high-coherence cavities pose increasing challenges for quantum control as coherence times improve. A central challenge lies in balancing fast, high-fidelity control and preserving the cavity’s exceptional coherence. Strong coupling to ancillary transmon circuits enables fast gates but introduces photon loss through the inverse Purcell effect and exposes the cavity to backaction from ancilla errors. One approach that resolves this tension is to reduce the coupling between the cavity and the lossy ancillary elements to protect cavity coherence, while overcoming the gate-speed limitations of weak coupling through techniques such as conditional displacements [2], sideband interactions [31], and sideband-enhanced selective number-dependent arbitrary phase (SNAP) gates [32]. Nevertheless, as coherence continues to improve, it becomes increasingly important to develop control strategies that are not only fast but also fault-tolerant to the dominant errors in the system [33].
@@
In this work, we present a two-mode bosonic quantum system implemented within a multimode SRF cavity, achieving ultrahigh coherence times and novel cavity control capabilities resilient to ancilla decoherence. Our system utilizes a two-cell niobium SRF cavity design based on the TESLA geometry [34, 35], weakly coupled to a transmon circuit for precise quantum control while preserving cavity coherence. We achieve single-photon lifetimes of 20.6 $\pm$ 0.4 ms and 15.6 $\pm$ 0.2 ms for the two bosonic modes, significantly longer than most state-of-the-art cavity QED platforms, establishing a record for multimode quantum memories. Despite the coherence-preserving weak coupling, we employ sideband operations to achieve fast cavity control on timescales far shorter than those set by the weak dispersive shift. In conjunction with novel protocols that make cavity control leading-order immune to ancilla errors, we achieve high-fidelity preparation of Fock states up to $N=20$ with fidelity exceeding 95%, as well as two-mode entanglement with estimated coherence-limited fidelities up to 99.9% after post-selection. These advancements in coherence and control mark a significant step in bosonic quantum computing, laying the groundwork for exploring high-dimensional qudit encodings and gate operations.
@@ II Cavity Characterization and System Design
We designed a two-mode superconducting cavity based on the multi-cell TESLA geometry [35, 36], consisting of two elliptically shaped cells connected by a circular aperture referred to as the iris, as shown in Fig. 1(a). The hybridization of the nearly degenerate bare TM010 modes in the two cells gives rise to two normal modes, named “Alice” and “Bob”, which have symmetric and antisymmetric electric field profiles, respectively (Fig. 1(b)). This normal mode structure leads to similar electric field density at the opening of the transmon tunnel, allowing for even coupling strength between the transmon and the two modes. The radius of the iris controls the coupling strength and consequently the frequency detuning between the normal modes. The Alice and Bob modes are designed to have frequencies of 5.779 GHz and 6.872 GHz, respectively. This frequency separation helps minimize crosstalk during control operations between the cavity modes, while also providing flexibility in placing the transmon frequency, which can vary due to fabrication inaccuracies. Cylindrical waveguide pipes with sufficiently high cut-off frequencies (35 GHz) are attached on either side of the cavity; one of these houses the transmon. Additionally, each pipe includes two ports for microwave drive lines to control transmon, readout and cavity modes.
@@
We manufactured this cavity from high-purity niobium, using e-beam welding to eliminate seam loss. State-of-the-art surface treatments, such as chemical etching and heat treatments, were applied to remove damaged surface layers and contaminants. The optimized surface quality and the intrinsic bulk property of high-purity niobium enable high quality factors for the cavity modes in the milli-Kelvin and low-photon-number regime [17, 37, 15, 16, 38], providing an excellent platform for quantum computing applications.
@@ III Ancilla-Error-Resilient Cavity State Preparation
An ancilla qubit dispersively coupled to a linear system enables universal control [39, 40, 29]. The resulting photon-number dependence of the ancilla transition frequency enables so-called SNAP gates, in which arbitrary phases can be applied to specific Fock states. Combined with cavity displacements, this allows for universal control of the cavity state [39, 1]. However, the standard SNAP gate duration should be sufficiently longer than $2\pi/\lvert\chi_{e}\rvert$, making it inefficient when $\chi_{e}$ is small. In contrast, both the Echoed Conditional Displacement (ECD) gate [41, 42, 43] and sideband control [44, 45, 31] also enable high-fidelity universal control, with a gate time controlled by microwave drive strength much shorter than the SNAP gate duration limit $2\pi/\lvert\chi_{e}\rvert$.
In this experiment, we utilize the sideband (SB) scheme [44, 40, 31] for controlling the cavity and further develop ancilla-error resilient state preparation schemes. The basic principle of preparing a large Fock state in a single mode is illustrated in Fig. 2(a). The transmon-resonator system is initially in state $|g,0\rangle$. The transmon is unconditionally excited to the $|f\rangle$ level applying broadband $\pi_{ge}$ and $\pi_{ef}$ pulses. Then a sideband $\pi_{\mathrm{sb}}^{(0)}$ pulse at the $|f,0\rangle\leftrightarrow|g,1\rangle$ transition is applied to prepare the state $|g1\rangle$ (see Appendix D for the calibration of the sideband pulse). By repeatedly applying a sequence of $\pi_{ge}$, $\pi_{ef}$, and $\pi_{\mathrm{sb}}^{(n)}$ sideband pulses at $|f,n\rangle\leftrightarrow|g,n+1\rangle$, it is possible to prepare Fock states up to a large photon number. It’s worth noting that the sideband pulses need not be frequency selective for Fock state preparation, or even for certain superposition states utilizing a recently developed shelving technique [31]. The sideband operation also allows for driven-dissipative reset of the cavity state in a time scale much faster than its natural lifetime (see Appendix E), which is utilized throughout our experiment.
@@
In climbing the Fock ladder to reach the target photon state, decoherence in both the transmon and cavity can introduce errors during state preparation. Because the cavity coherence time significantly exceeds that of the transmon, and since transmon $\pi$ pulses in our system are approximately twenty times faster than sideband operations, the fidelity is primarily limited by transmon errors occurring during sideband pulses. Specifically, a transmon decay event from the initial state $|f,n\rangle$ results in the error state $|e,n\rangle$, preventing the sideband transition to $|g,n+1\rangle$; similarly, transmon dephasing disrupts the coherent transition $|f,n\rangle$ to $|g,n+1\rangle$, leaving residual population in $|f,n\rangle$. Although the probability of such events within a single sideband operation is low, the multiple repetitions required to prepare high-photon-number Fock states can cumulatively lead to significant infidelity. To mitigate this cumulative infidelity caused by the ancilla, we have developed a “sideband feedforward protocol” (SFP), illustrated in Fig. 2(b). Akin to other techniques exploiting multi-level transmon readout for fault-tolerant cavity operations [46, 47, 48, 49, 50], the SFP leverages error syndrome detection via high-fidelity single-shot measurements of the transmon populations in the $|g\rangle$, $|e\rangle$, and $|f\rangle$ states after each sideband operation. This information is then fed forward to determine the subsequent corrective pulse in real time. A measurement outcome of $|f\rangle$ heralds a dephasing event, which we correct by repeating the same sideband pulse. An outcome of $|e\rangle$ indicates a transmon decay event; in this case, we first apply a $\pi_{ef}$ pulse and then repeat the sideband pulse to restore the intended state. Finally, an outcome of $|g\rangle$ indicates the ideal scenario where no error has occurred, and thus requires no further correction.
@@
We plot the state preparation fidelities of Fock states up to twenty photons for Alice and Bob modes in Fig. 2(d). The traces show clear improvement of the Fock state fidelity obtained with SFP and PF techniques, compared to the vanilla SB protocol. We achieve preparation fidelities of 95.3±1.9%95.3\ $\pm$ 1.9\ $\%$ and 94.6± 1.1%94.6\ $\pm\$$1\%$ for the Fock state $|\psi\rangle_{A/B}=|20\rangle$ in the Alice and Bob modes, respectively. To the best of our knowledge, this represents the highest reported fidelities for such high-photon-number Fock states to date. The high-fidelity Fock state preparation also enables calibration of sideband control at even higher photon numbers, preserving high contrast in the calibration measurement. While we have demonstrated only basis state preparation, sideband control enables arbitrary state synthesis [40, 51] and theoretically supports universal multi-qudit gate operations [52, 45]. Achieving high fidelity and long coherence for large Fock states thus opens new opportunities for qudit-based quantum computation [22, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63], quantum sensing [64, 65], and quantum simulation [66, 67, 28, 68, 30, 69, 70, 29, 71].
@@ IV Sideband-mediated Virtual Raman Interaction for Cavity Entanglement
As a demonstration of two-mode control in this cavity system, we extend the sideband approach to implement a two-mode entangling operation, realizing a beamsplitter interaction in the single-photon subspace of the two modes. Such an interaction has previously been implemented using either a transmon [72, 73] or a dedicated coupling element to enable three- or four-wave mixing between modes [6, 7]. In contrast to standard four-wave mixing with a fixed-frequency transmon, we leverage the transmon’s $|f\rangle$ level to mediate the interaction. By simultaneously driving the transmon at a detuning $\Delta$ from the $|f0\rangle\leftrightarrow|g1\rangle$ sideband transitions of both Alice and Bob, we activate a virtual-Raman-assisted beamsplitter (VRBS) interaction between the two modes, enabling coherent photon exchange and entanglement (see Fig. 3(a)). Choosing $\Delta$ larger than the sideband rates suppresses direct population of the more lossy transmon $|f\rangle$ state during the interaction, albeit at the cost of a reduced conversion rate. Importantly, compared to the standard four-wave mixing protocol, our technique achieves a faster conversion rate for a given drive strength, with an enhancement factor of $|\alpha/\Delta|$, where $\alpha$ is the transmon anharmonicity (see Appendix I).
To estimate the fidelity of the entangling operation, we measure the evolution of the two modes under the VRBS interaction, with an initial single-photon Fock state in the Alice mode. After an evolution time of $t$, we turn off the VRBS interaction, and map the two-mode states in the single-photon subspace, specifically $\rangle, |01|01\rangle$, and the erasure state $|00\rangle$ (order: Alice $\otimes$ Bob), onto the transmon states $|e\rangle$, $|f\rangle$, and $|g\rangle$, respectively (see Appendix M). This allows us to monitor the evolution of the state populations as a function of time, which can be fitted to the theoretical model described in [7],
@@
We optimize these fidelities as a function of the VRBS detuning $\Delta$, under fixed sideband drive strengths, $g_{sb}^{a}/2\pi=269.5\pm 0.5$ kHz, $g_{sb}^{b}/2\pi=301.8\pm 0.5$ kHz. The Raman-assisted beamsplitter rate increases as $\Delta$ decreases, as shown by the experimental data in Fig. 3(b, left). However, this enhanced interaction comes at the cost of both increased decay and dephasing. Qualitatively, the beamsplitter rate scales inversely with $\Delta$ (see Appendix I), while the sideband-drive-induced Purcell decay follows an approximate $\Delta^{-2}$ dependence [32]. This scaling alone would suggest that the total infidelity should decrease with increasing detuning. However, at large detunings, cavity decoherence is no longer limited by sideband-induced decay, but is instead dominated by dephasing due to transmon heating induced by the strong drive tones. This heating arises from undesired parametric processes such as dressed dephasing [74]. A heating event excites the transmon to the $\lvert e\rangle$ state, effectively halting the coherent photon exchange. These random transmon heating events dephase the VRBS oscillation as a result of the strong transmon-state-dependence of the beamsplitter rate (see Appendix I). Additionally, transmon heating enhances photon shot noise dephasing of the cavity modes. Since these heating processes are governed by the detuning between the transmon and the drive (on the order of GHz) [75], they are largely insensitive to MHz-scale variations in $\Delta$, leading to the observed saturation of the beamsplitter dephasing time $1/\kappa_{\phi}$ at large detunings in Fig. 3(b, middle).
@@
The coherent photon exchange process corresponding to the VRBS interaction at the optimal parameters are shown in Fig. 3(c, top). Fitting this oscillation using Eq. (2) yields a coherence-limited beamsplitter fidelity of $99.686\pm 0.004\%$. This operation can be interpreted as a single-qubit gate on a dual-rail qubit ($\{|10\rangle,|01\rangle\}$) encoded across the single-photon subspace of the two cavity modes [50, 76]. We observe a gradual increase in the population of the $|00\rangle$ state during the interaction as a result of photon loss in the cavity modes, equivalent to the erasure error of a dual-rail qubit. While we do not implement an erasure syndrome measurement that preserves the dual-rail qubit state, we estimate the improvement in fidelity from erasure checks by post-selecting on the absence of erasures at the end of the oscillation. The resulting oscillation, conditioned on no erasure events, exhibits a prolonged decay constant of $1/\kappa_{1}=70.85$ ms, translating to an improved $\mathcal{F}^{*}_{\mathrm{BS}}$ of $99.810\pm 0.003\%$ (Fig. 3(c, bottom)).
@@
In addition to post-selecting on erasure errors, we also perform mid-circuit detection of transmon heating errors, which, as previously discussed, are the dominant source of dephasing in the VRBS operation. These heating errors are detected via a transmon measurement, which—when post-selecting on the transmon being in the ground state—introduces negligible backaction on the cavity states due to the weak hybridization between the transmon and the cavities in our regime. We verify that the cavity Fock state lifetimes are unaffected by repeated transmon measurements in Appendix M. We demonstrate the resulting fidelity improvement using a SWAP operation. To amplify the effect of heating, we apply a sequence of SWAP gates implemented via the same VRBS interaction. As a baseline, Fig. 3(d, top) shows the outcome of this sequence—initialized in $\lvert g10\rangle$—with only a final erasure check. The measured heating rate is found to be approximately 1% per SWAP (See Appendix L). To quantify the benefit of heating error detection, we then perform transmon measurements after each SWAP. The corresponding results, also shown in Fig. 3(d, middle), exhibit a clear improvement in contrast after one hundred SWAPs, corresponding to a two-fold improvement in fidelity (infidelity drops from $0.4\%\rightarrow 0.2\%$). We further verify that this improvement holds for initial states prepared as entangled superpositions of the two modes (i.e., equatorial states on the Bloch sphere in the dual-rail encoding) by applying a $U_{\mathrm{BS}}$ following the initialization of a Fock state in one mode (Fig. 3(d, bottom)), and another $U_{\mathrm{BS}}(\phi)$ before the readout to account for Stark-shift-induced phase rotations (see Appendix K). We thereby realize a partially error-detected SWAP gate, which can be incorporated mid-circuit to improve the fidelity of shallow-depth algorithms and to support dual-rail surface code implementations [49] using simpler fixed-frequency transmon couplers. This error-detection scheme can also be extended to beamsplitter operations via additional pulse calibration steps that compensate for Stark shift effects [7]. Based on the SWAP gate measurement results, we infer a coherence-limited fidelity approaching $99.9\%$ for the error-detected beamsplitter gate—comparable to state-of-the-art results, albeit with longer gate durations, but achieved using significantly simpler circuit hardware.

### v3

@@ Abstract
Bosonic quantum systems have emerged as a compelling platform for quantum computing, offering several intrinsic advantages over traditional qubit-based architectures. By encoding quantum information in harmonic oscillator modes, these systems benefit from long coherence times, large Hilbert spaces, and relatively simple error channels dominated by photon loss. Coupling to nonlinear superconducting circuits has enabled universal control and gate operations using a variety of techniques [SNAP2015PRL, Eickbusch2022, rosenblum2018cnot, heeres2017implementing, gao2019entanglement, PRXQuantum.4.020355, LuSchoelkopf2023, Maiti2025] in both single-mode [reagor2016quantum] and multimode [wang2016schrodinger, Chakram2021seamless] circuit QED systems implemented with three-dimensional (3D) SRF cavities featuring millisecond-scale coherence times. These capabilities have enabled hardware-efficient implementations of bosonic quantum error correction, including demonstrations that surpass the break-even point [ofek2016extending, sivak2023real, ni2023beating]. More recently, these control strategies have begun to extend to ultra-high-coherence cavities [Rosenblum2023mushroom, Oriani2024Nb_coax], with coherence times exceeding tens of milliseconds and the potential for further improvements [Romanenko2020], opening new avenues for scalable quantum information processing.
@@
Qubit-based quantum computing has long been the dominant paradigm. In bosonic error correction, cavity modes are typically used to encode logical qubits [mirrahimi2014dynamically, michael2016new]. The large Hilbert space of a cavity also makes it a natural platform for implementing a qudit using multiple quantum levels, both natively and with error correction [brock2025quantum]. Qudits offer enhanced information density and can enable more compact and efficient quantum circuits across a wide range of algorithms [bullock2005asymptotically, wang2020qudits, low2025]. They are particularly well-suited for quantum simulations in chemistry [cao2019quantum, dutta2024simulating], condensed matter [Chizzini2024], and high-energy physics (HEP) [Daniel2022, Zache:2023cfj], where they can represent physical degrees of freedom such as large spins, rotors, and gauge fields [Girvin2024hybrid]. In HEP applications, a quantum field can also be discretized in position space and encoded within a single high-dimensional qudit [gustafson2022primitive]. The emergence of high-coherence cavities presents a promising path for improved control over high-photon-number Fock states, thereby enabling qudit-based quantum computing and simulation.
@@
Despite their advantages, high-coherence cavities pose increasing challenges for quantum control as coherence times improve. A central challenge lies in balancing fast, high-fidelity control and preserving the cavity’s exceptional coherence. Strong coupling to ancillary transmon circuits enables fast gates but introduces photon loss through the inverse Purcell effect and exposes the cavity to backaction from ancilla errors. One approach that resolves this tension is to reduce the coupling between the cavity and the lossy ancillary elements to protect cavity coherence, while overcoming the gate-speed limitations of weak coupling through techniques such as conditional displacements [Eickbusch2022], sideband interactions [huang2025fast], and sideband-enhanced selective number-dependent arbitrary phase (SNAP) gates [you2025floquet]. Nevertheless, as coherence continues to improve, it becomes increasingly important to develop control strategies that are not only fast but also fault-tolerant to the dominant errors in the system [pietikainen2024strategies].
@@
In this work, we present a two-mode bosonic quantum system implemented within a multimode SRF cavity, achieving ultrahigh coherence times and novel cavity control capabilities resilient to ancilla decoherence. Our system utilizes a two-cell niobium SRF cavity design based on the TESLA geometry [Padamsee1998book, Aune2000], weakly coupled to a transmon circuit for precise quantum control while preserving cavity coherence. We achieve single-photon lifetimes of 20.6 $\pm$ 0.4 ms and 15.6 $\pm$ 0.2 ms for the two bosonic modes, significantly longer than most state-of-the-art cavity QED platforms, establishing a record for multimode quantum memories. Despite the coherence-preserving weak coupling, we employ sideband operations to achieve fast cavity control on timescales far shorter than those set by the weak dispersive shift. In conjunction with novel protocols that make cavity control leading-order immune to ancilla errors, we achieve high-fidelity preparation of Fock states up to $N=20$ with fidelity exceeding 95%, as well as two-mode entanglement with estimated coherence-limited fidelities up to 99.9% after post-selection. These advancements in coherence and control mark a significant step in bosonic quantum computing, laying the groundwork for exploring high-dimensional qudit encodings and gate operations.
@@ II Cavity Characterization and System Design
We designed a two-mode superconducting cavity based on the multi-cell TESLA geometry [Aune2000, Reineri2023], consisting of two elliptically shaped cells connected by a circular aperture referred to as the iris, as shown in Fig. 1(a). The hybridization of the nearly degenerate bare TM010 modes in the two cells gives rise to two normal modes, named “Alice” and “Bob”, which have symmetric and antisymmetric electric field profiles, respectively (Fig. 1(b)). This normal mode structure leads to similar electric field density at the opening of the transmon tunnel, allowing for even coupling strength between the transmon and the two modes. The radius of the iris controls the coupling strength and consequently the frequency detuning between the normal modes. The Alice and Bob modes are designed to have frequencies of 5.779 GHz and 6.872 GHz, respectively. This frequency separation helps minimize crosstalk during control operations between the cavity modes, while also providing flexibility in placing the transmon frequency, which can vary due to fabrication inaccuracies. Cylindrical waveguide pipes with sufficiently high cut-off frequencies (35 GHz) are attached on either side of the cavity; one of these houses the transmon. Additionally, each pipe includes two ports for microwave drive lines to control transmon, readout and cavity modes.
@@
We manufactured this cavity from high-purity niobium, using e-beam welding to eliminate seam loss. State-of-the-art surface treatments, such as chemical etching and heat treatments, were applied to remove damaged surface layers and contaminants. The optimized surface quality and the intrinsic bulk property of high-purity niobium enable high quality factors for the cavity modes in the milli-Kelvin and low-photon-number regime [Romanenko2020, Paul2021, Rosenblum2023mushroom, Oriani2024Nb_coax, charpentier2025], providing an excellent platform for quantum computing applications.
@@ III Ancilla-Error-Resilient Cavity State Preparation
An ancilla qubit dispersively coupled to a linear system enables universal control [SNAP2015PRA, law1996arbitrary, Girvin2024hybrid]. The resulting photon-number dependence of the ancilla transition frequency enables so-called SNAP gates, in which arbitrary phases can be applied to specific Fock states. Combined with cavity displacements, this allows for universal control of the cavity state [SNAP2015PRA, SNAP2015PRL]. However, the standard SNAP gate duration should be sufficiently longer than $2\pi/\lvert\chi_{e}\rvert$, making it inefficient when $\chi_{e}$ is small. In contrast, both the Echoed Conditional Displacement (ECD) gate [ECD2022NatPhys, Diringer2024CNOD, You2024crosstalk] and sideband control [sideband2015PRA, Liu2021qudit, huang2025fast] also enable high-fidelity universal control, with a gate time controlled by microwave drive strength much shorter than the SNAP gate duration limit $2\pi/\lvert\chi_{e}\rvert$.
In this experiment, we utilize the sideband (SB) scheme [sideband2015PRA, law1996arbitrary, huang2025fast] for controlling the cavity and further develop ancilla-error resilient state preparation schemes. The basic principle of preparing a large Fock state in a single mode is illustrated in Fig. 2(a). The transmon-resonator system is initially in state $|g,0\rangle$. The transmon is unconditionally excited to the $|f\rangle$ level applying broadband $\pi_{ge}$ and $\pi_{ef}$ pulses. Then a sideband $\pi_{\mathrm{sb}}^{(0)}$ pulse at the $|f,0\rangle\leftrightarrow|g,1\rangle$ transition is applied to prepare the state $|g1\rangle$ (see Appendix D for the calibration of the sideband pulse). By repeatedly applying a sequence of $\pi_{ge}$, $\pi_{ef}$, and $\pi_{\mathrm{sb}}^{(n)}$ sideband pulses at $|f,n\rangle\leftrightarrow|g,n+1\rangle$, it is possible to prepare Fock states up to a large photon number. It’s worth noting that the sideband pulses need not be frequency selective for Fock state preparation, or even for certain superposition states utilizing a recently developed shelving technique [huang2025fast]. The sideband operation also allows for driven-dissipative reset of the cavity state in a time scale much faster than its natural lifetime (see Appendix E), which is utilized throughout our experiment.
@@
In climbing the Fock ladder to reach the target photon state, decoherence in both the transmon and cavity can introduce errors during state preparation. Because the cavity coherence time significantly exceeds that of the transmon, and since transmon $\pi$ pulses in our system are approximately twenty times faster than sideband operations, the fidelity is primarily limited by transmon errors occurring during sideband pulses. Specifically, a transmon decay event from the initial state $|f,n\rangle$ results in the error state $|e,n\rangle$, preventing the sideband transition to $|g,n+1\rangle$; similarly, transmon dephasing disrupts the coherent transition $|f,n\rangle$ to $|g,n+1\rangle$, leaving residual population in $|f,n\rangle$. Although the probability of such events within a single sideband operation is low, the multiple repetitions required to prepare high-photon-number Fock states can cumulatively lead to significant infidelity. To mitigate this cumulative infidelity caused by the ancilla, we have developed a “sideband feedforward protocol” (SFP), illustrated in Fig. 2(b). Akin to other techniques exploiting multi-level transmon readout for fault-tolerant cavity operations [Rosenblum2018, Elder2020, Reinhold2020, Taka2023, teoh2023dual], the SFP leverages error syndrome detection via high-fidelity single-shot measurements of the transmon populations in the $|g\rangle$, $|e\rangle$, and $|f\rangle$ states after each sideband operation. This information is then fed forward to determine the subsequent corrective pulse in real time. A measurement outcome of $|f\rangle$ heralds a dephasing event, which we correct by repeating the same sideband pulse. An outcome of $|e\rangle$ indicates a transmon decay event; in this case, we first apply a $\pi_{ef}$ pulse and then repeat the sideband pulse to restore the intended state. Finally, an outcome of $|g\rangle$ indicates the ideal scenario where no error has occurred, and thus requires no further correction.
@@
We plot the state preparation fidelities of Fock states up to twenty photons for Alice and Bob modes in Fig. 2(d). The traces show clear improvement of the Fock state fidelity obtained with SFP and PF techniques, compared to the vanilla SB protocol. We achieve preparation fidelities of 95.3±1.9%95.3\ $\pm$ 1.9\ $\%$ and 94.6± 1.1%94.6\ $\pm\$$1\%$ for the Fock state $|\psi\rangle_{A/B}=|20\rangle$ in the Alice and Bob modes, respectively. To the best of our knowledge, this represents the highest reported fidelities for such high-photon-number Fock states to date. The high-fidelity Fock state preparation also enables calibration of sideband control at even higher photon numbers, preserving high contrast in the calibration measurement. While we have demonstrated only basis state preparation, sideband control enables arbitrary state synthesis [law1996arbitrary, hofheinz2009synthesizing] and theoretically supports universal multi-qudit gate operations [mischuck2013qudit, Liu2021qudit]. Achieving high fidelity and long coherence for large Fock states thus opens new opportunities for qudit-based quantum computation [wang2020qudits, Blok2021, Goss2022, Roy2023two_qutrit, nguyen2024empowering, Nikolaeva:2024wxl, Gustafson:2025buk, Litteken2023quantum_waltz, Cao2024transmon4, Wang2025transmon12, Blok2025transmon8, shi2025Grover_qudit], quantum sensing [Agrawal2024stimulated, deng2024metrology], and quantum simulation [gustafson2021prospects, ciavarella2021trailhead, Zache:2023cfj, bauer2023quantum, gustafson2022primitive, illa2023quantum, Kurkcuoglu:2024cfv, Girvin2024hybrid, meth2025LG_qudit].
@@ IV Sideband-mediated Virtual Raman Interaction for Cavity Entanglement
As a demonstration of two-mode control in this cavity system, we extend the sideband approach to implement a two-mode entangling operation, realizing a beamsplitter interaction in the single-photon subspace of the two modes. Such an interaction has previously been implemented using either a transmon [Gao2018, pfaff2017controlled] or a dedicated coupling element to enable three- or four-wave mixing between modes [PRXQuantum.4.020355, LuSchoelkopf2023]. In contrast to standard four-wave mixing with a fixed-frequency transmon, we leverage the transmon’s $|f\rangle$ level to mediate the interaction. By simultaneously driving the transmon at a detuning $\Delta$ from the $|f0\rangle\leftrightarrow|g1\rangle$ sideband transitions of both Alice and Bob, we activate a virtual-Raman-assisted beamsplitter (VRBS) interaction between the two modes, enabling coherent photon exchange and entanglement (see Fig. 3(a)). Choosing $\Delta$ larger than the sideband rates suppresses direct population of the more lossy transmon $|f\rangle$ state during the interaction, albeit at the cost of a reduced conversion rate. Importantly, compared to the standard four-wave mixing protocol, our technique achieves a faster conversion rate for a given drive strength, with an enhancement factor of $|\alpha/\Delta|$, where $\alpha$ is the transmon anharmonicity (see Appendix I).
To estimate the fidelity of the entangling operation, we measure the evolution of the two modes under the VRBS interaction, with an initial single-photon Fock state in the Alice mode. After an evolution time of $t$, we turn off the VRBS interaction, and map the two-mode states in the single-photon subspace, specifically $\rangle, |01|01\rangle$, and the erasure state $|00\rangle$ (order: Alice $\otimes$ Bob), onto the transmon states $|e\rangle$, $|f\rangle$, and $|g\rangle$, respectively (see Appendix M). This allows us to monitor the evolution of the state populations as a function of time, which can be fitted to the theoretical model described in [LuSchoelkopf2023],
@@
We optimize these fidelities as a function of the VRBS detuning $\Delta$, under fixed sideband drive strengths, $g_{sb}^{a}/2\pi=269.5\pm 0.5$ kHz, $g_{sb}^{b}/2\pi=301.8\pm 0.5$ kHz. The Raman-assisted beamsplitter rate increases as $\Delta$ decreases, as shown by the experimental data in Fig. 3(b, left). However, this enhanced interaction comes at the cost of both increased decay and dephasing. Qualitatively, the beamsplitter rate scales inversely with $\Delta$ (see Appendix I), while the sideband-drive-induced Purcell decay follows an approximate $\Delta^{-2}$ dependence [you2025floquet]. This scaling alone would suggest that the total infidelity should decrease with increasing detuning. However, at large detunings, cavity decoherence is no longer limited by sideband-induced decay, but is instead dominated by dephasing due to transmon heating induced by the strong drive tones. This heating arises from undesired parametric processes such as dressed dephasing [Boissonneault2008]. A heating event excites the transmon to the $\lvert e\rangle$ state, effectively halting the coherent photon exchange. These random transmon heating events dephase the VRBS oscillation as a result of the strong transmon-state-dependence of the beamsplitter rate (see Appendix I). Additionally, transmon heating enhances photon shot noise dephasing of the cavity modes. Since these heating processes are governed by the detuning between the transmon and the drive (on the order of GHz) [yaxing], they are largely insensitive to MHz-scale variations in $\Delta$, leading to the observed saturation of the beamsplitter dephasing time $1/\kappa_{\phi}$ at large detunings in Fig. 3(b, middle).
@@
The coherent photon exchange process corresponding to the VRBS interaction at the optimal parameters are shown in Fig. 3(c, top). Fitting this oscillation using Eq. (2) yields a coherence-limited beamsplitter fidelity of $99.686\pm 0.004\%$. This operation can be interpreted as a single-qubit gate on a dual-rail qubit ($\{|10\rangle,|01\rangle\}$) encoded across the single-photon subspace of the two cavity modes [teoh2023dual, chou2024superconducting]. We observe a gradual increase in the population of the $|00\rangle$ state during the interaction as a result of photon loss in the cavity modes, equivalent to the erasure error of a dual-rail qubit. While we do not implement an erasure syndrome measurement that preserves the dual-rail qubit state, we estimate the improvement in fidelity from erasure checks by post-selecting on the absence of erasures at the end of the oscillation. The resulting oscillation, conditioned on no erasure events, exhibits a prolonged decay constant of $1/\kappa_{1}=70.85$ ms, translating to an improved $\mathcal{F}^{*}_{\mathrm{BS}}$ of $99.810\pm 0.003\%$ (Fig. 3(c, bottom)).
@@
In addition to post-selecting on erasure errors, we also perform mid-circuit detection of transmon heating errors, which, as previously discussed, are the dominant source of dephasing in the VRBS operation. These heating errors are detected via a transmon measurement, which—when post-selecting on the transmon being in the ground state—introduces negligible backaction on the cavity states due to the weak hybridization between the transmon and the cavities in our regime. We verify that the cavity Fock state lifetimes are unaffected by repeated transmon measurements in Appendix M. We demonstrate the resulting fidelity improvement using a SWAP operation. To amplify the effect of heating, we apply a sequence of SWAP gates implemented via the same VRBS interaction. As a baseline, Fig. 3(d, top) shows the outcome of this sequence—initialized in $\lvert g10\rangle$—with only a final erasure check. The measured heating rate is found to be approximately 1% per SWAP (See Appendix L). To quantify the benefit of heating error detection, we then perform transmon measurements after each SWAP. The corresponding results, also shown in Fig. 3(d, middle), exhibit a clear improvement in contrast after one hundred SWAPs, corresponding to a two-fold improvement in fidelity (infidelity drops from $0.4\%\rightarrow 0.2\%$). We further verify that this improvement holds for initial states prepared as entangled superpositions of the two modes (i.e., equatorial states on the Bloch sphere in the dual-rail encoding) by applying a $U_{\mathrm{BS}}$ following the initialization of a Fock state in one mode (Fig. 3(d, bottom)), and another $U_{\mathrm{BS}}(\phi)$ before the readout to account for Stark-shift-induced phase rotations (see Appendix K). We thereby realize a partially error-detected SWAP gate, which can be incorporated mid-circuit to improve the fidelity of shallow-depth algorithms and to support dual-rail surface code implementations [Taka2023] using simpler fixed-frequency transmon couplers. This error-detection scheme can also be extended to beamsplitter operations via additional pulse calibration steps that compensate for Stark shift effects [LuSchoelkopf2023]. Based on the SWAP gate measurement results, we infer a coherence-limited fidelity approaching $99.9\%$ for the error-detected beamsplitter gate—comparable to state-of-the-art results, albeit with longer gate durations, but achieved using significantly simpler circuit hardware.